# Accepted with Minor Revisions: Value of AI-Assisted Scientific Writing

### v1

@@ Abstract.
Large Language Models have seen expanding application across domains, yet their effectiveness as assistive tools for scientific writing—an endeavor requiring precision, multimodal synthesis, and domain expertise—remains insufficiently understood. We examine the potential of LLMs to support domain experts in scientific writing, with a focus on abstract composition. We design an incentivized randomized controlled trial with a hypothetical conference setup where participants with relevant expertise are split into an author and reviewer pool. Inspired by methods in behavioral science, our novel incentive structure encourages authors to edit the provided abstracts to an acceptable quality for a peer-reviewed submission. Our $2\times 2$ between-subject design expands into two dimensions: the implicit source of the provided abstract and the disclosure of it. We find authors make most edits when editing human-written abstracts compared to AI-generated abstracts without source attribution, often guided by higher perceived readability in AI generation. Upon disclosure of source information, the volume of edits converges in both source treatments. Reviewer decisions remain unaffected by the source of the abstract, but bear a significant correlation with the number of edits made. Careful stylistic edits, especially in the case of AI-generated abstracts, in the presence of source information, improve the chance of acceptance. We find that AI-generated abstracts hold potential to reach comparable levels of acceptability to human-written ones with minimal revision, and that perceptions of AI authorship, rather than objective quality, drive much of the observed editing behavior. Our findings reverberate the significance of source disclosure in collaborative scientific writing.
@@
$c$s: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Telling people about research is as important as doing it. Scientific publishing is an integral part of science, meant to disseminate research findings, foster collaboration, encourage reproducibility, and ensure that scientific knowledge is accessible and built upon over time (liang2024mapping). Bad writing can and often does prevent or delay the publication of good science (gastel2022write). Researchers and scientists today, however, are often not formally trained to write, and generally learn by imitating the styles of their advisors and other authors (lindsay2020scientific). Large language models (LLMs), with their writing abilities, offer a promising venue to assist researchers in both conducting and writing about research. However, a thorough assessment of their capabilities in scientific writing is lacking.
@@
Even so, the use of language models continues to grow in this area. Many tools and platforms termed “AI co-scientist” have emerged in the past year (gottweis2025towards; Beel2025EvaluatingSA). They are marketed to be capable of generating, executing, and writing papers, with some of them getting accepted in academic workshops organized by peer-reviewed conferences (yamada2025aiscientistv2workshoplevelautomated). Popular scientific writing platforms such as Overleaf now offer writing assistant plugins. [^1] https://www.overleaf.com/learn/how-to/Writefull_integration However, anecdotal examples of both seemingly AI-generated papers (cybernews_academic_2023; anthes_ai_2023) and peer reviews (oransky_papers_2024) have inspired criticisms (liang2024mapping). In addition, across disciplines, recent studies have raised concerns over the declining quality of writing in scientific publications in several disciplines, attributing it to the rise of LLM-generated content (matsui2024delving; lemire2024will; geng2024chatgpt). These studies, however, have largely focused on corpus-level trends on unreviewed manuscripts and are conducted without consulting or involving writers such as the authors of said publications.
To raise awareness and prevent misuse by bad actors or unintentional use by over-reliant writers, there is an urgent need to conduct a systematic evaluation of the writing abilities of the current state-of-the-art models in specialized domains. Furthermore, despite current limitations, AI assistance holds tremendous potential in accelerating future scientific progress (Sourati2023AcceleratingSW; Nathani2025MLGymAN; Majumder2024DatadrivenDW). Towards understanding scientific writing capabilities, we seek to answer the following research question in this work: maintaining the standard for scientific writing quality, to what extent do researchers rely on model-generated content for academic publishing?
@@
We simulate an abstract editing and reviewing process mimicking the process of scientific publication in a laboratory setting using the methodology of incentivized behavioral experiments (azrieli2018incentives) to encourage participants to perform the task to the best of their ability. Authors edit the abstracts, and we quantify author reliance based on the extent of their modifications to make them submission-ready for a hypothetical conference publication. The task is conducted under two information conditions: (1) authors are informed that the original abstract provided to them was AI-generated, and (2) authors are not informed that the abstract was assembled by generative AI. To contextualize this behavior with a more realistic workflow, we also provide the authors with human-written abstracts under the same information conditions. By considering abstracts from different sources (human experts and LLMs), we assess the reception of LLM-generated abstracts by evaluating correlations between acceptance rates, edit frequency, and edit types.
@@
The edited abstracts are then evaluated by a group of (incentivized) reviewers in a double-blind fashion, who vote to accept or reject the submitted abstract. This design allows us to examine how authors’ reliance varies across source types and information conditions, and how expert reviewers perceive the quality and credibility of abstracts produced through generative AI workflows. These experiments result in a collection of model-generated and human-edited texts, where we capture keystroke-level edits. From this rich data of inputs, abstracts, and edits, we conduct several evaluations and analyses. Our main evaluation is based on the correlation between the rate of editing and the final rating an abstract receives, contrasting the controls and different treatments. We consider and measure the impact of variables like writer expertise (undergraduate students, graduate students, experienced researchers), and demographics and social factors (gender, perception of AI). In addition, we quantitatively (using linguistic tools) and qualitatively (via interviews with participants) characterize different kinds of edits authors make to understand the differences between edits made to AI- and human-authored abstracts.
@@
Across our experiments, we find that both the implicit origin of abstract and the disclosure of it systematically shape editing behavior and outcomes. Without source disclosure, authors make fewer edits to AI-generated abstracts, perceiving them as more readable, though PhD-level authors edited AI text more extensively. Despite these differences, reviewer evaluations showed no differences in accepting abstracts edited under no disclosure conditions. Disclosure of source primarily triggers social adjustments: authors reduced edits for human-written text when the source was revealed, whereas disclosure had only minor effects on AI-generated abstracts, with small impacts on perceived readability and confidence. Our quantitative analysis of stylistic metrics reveals that edits to AI-generated abstracts are to improve cohesion, reduce nominalizations, and produce informative sentences. On the contrary, authors tend to lengthen the openings of the human-written abstracts. Similarly, disclosure of source information alone has a minimal direct effect on the writing style. Finally, thematic analysis of our interviews with authors confirms that authors adopt different strategies depending on the origin of the abstracts: they simplify AI-generated text, restructure human-authored abstracts for emphasis, and disclosure promotes the sense of accountability equally for human and AI. Together, through comprehensive large-scale experiments, we first highlight [^2] Our experimental data and instructions are provided as supplementary. the varying cognitive editorial strategies emerging from writing sessions that involve both human and AI-generated scientific writing.
@@ 2.1. AI Systems for Automating Scientific Research
Early works on AI-assistance in scientific research largely focused on surface-level corrections and sentence-level improvements in writing (Ito2019DiamondsIT; Bezerra2021NEWRITERAT) with some attempts at drafting entire paper structure (wang-etal-2019-paperrobot). Recent work has started to apply LLMs to the entire research cycle, including hypothesis discovery (zhou2024hypothesis; zhang2025exploring; majumder2024discoverybench; abdel2025scientific), experiment planning and execution (li2025can; binbas2025mobllm; gui2025leveraging), scientific writing (kanna2024much; Liebling2025TowardsAA; Kousha2025HowMA), and writing reviews (silva2025ai). Within the writing stage, some works have focused on specific subtasks like metaphor generation (kim2023metaphorian), figure caption generation (kim2025multi), citation generation (xing-etal-2020-automatic; fierro2024learning; jung2022intent), and related work generation (https://doi.org/10.48550/arxiv.2212.09577; li2024related). Such tasks are, typically, accompanied by reference answers, making evaluation relatively straightforward. More recent work has started focusing on generating entire papers without any human intervention (lu2024ai). In this work, we adopt a more realistic setup of humans relying on an AI-assistant to generate drafts, which they then edit (kanna2024much). We also use a principled method of human evaluation in the form peer review conducted via incentivized experiments.
A wide range of commercial [^3] https://www.gatsbi.com/, https://anara.com/blog/ai-research-tools as well as open-source tools now offer services for every stage of the writing process, from literature review and synthesis to full manuscript drafting, editing, and refinement (shao2025sciscigpt). Specific features, such as automated abstract and title generators, are also becoming commonplace [^4] https://www.writefull.com/. The widespread availability and marketing of these tools signify a de facto adoption that necessitates rigorous empirical evaluation of their effectiveness, their impact on the quality of scientific output, and the human behaviors they engender, which our study addresses.
@@ 2.2. Human AI Collaboration and Co-writing
The majority of prior research on Human-AI co-writing has focused on creative writing (singh2025systematic). Using LLMs, researchers have developed and evaluated systems supporting story writing (yuan2022wordcraft), playwriting (10.1145/3656650.3656688), and character development (park2025constella). Some works also explore higher-level writing tasks such as prewriting (wan2024felt), and generating perspective-specific feedback (rashkin2025help). Specialized creative applications have emerged for tasks including metaphor generation (van2024perspective), collaborative storytelling (10.1145/3649921.3656987), and personal diary writing (10.1145/3613904.3642693). User studies reveal complex dynamics in how writers interact with and perceive AI writing assistance. At a system interaction level, the design of AI suggestions significantly impacts user behavior and output: sentence-level suggestions promote original content creation, while paragraph-level suggestions improve efficiency (10.1145/3544548.3581351; 10.1145/3613904.3642134). Writers’ engagement with AI assistance is also influenced by their personal values and goals (10.1145/3532106.3533506; guo2025pen). Writers show varying receptivity to AI support based on their confidence levels, demonstrating higher acceptance in areas where they lack expertise (al2025adoption; wang2025survey), and their desires for support are closely tied to their perception of support actors and personal values (hwang202580; guo2024preserving). Professional writers note persistent challenges with AI systems’ ability to maintain consistent style and voice (chakrabarty2024art).
Moreover, this Human-AI writing relationship raises important concerns. Studies reveal that biased AI models can influence not only the resulting text but also users’ own opinions (jakesch2023co). Prior research has identified biases, quality issues, lack of structure, and superficiality in AI-generated text, especially in domains like fiction and satire (chakrabarty2024art; chakrabarty2024creativity; ippolito2022creative; marco2024pron; mirowski2023co; mirowski2024robot). Researchers have also found LLM-generated content to be overly positive and lacking in nuance and have highlighted the need for more diverse and representative training data to mitigate these shortcomings (tian2024large). Our work adapts evaluation practices from prior studies in co-writing by focusing on the quality of AI-generated text using the amount of edits made to it by a human co-writer.
@@ 2.3. Cognitive Biases in Human-AI interaction
The inability of humans to reliably detect AI-generated text makes their beliefs about a text’s origin a powerful psychological variable, opening the door for cognitive biases to influence their judgment and behavior (Jakesch2022HumanHF; henestrosa2024understanding; FIEDLER2025100321). A rich literature in HCI and psychology documents two opposing, context-dependent human tendencies when interacting with automated systems (844354; lee2004trust; DZINDOLET2003697). On one hand, automation bias describes the tendency to over-rely on or excessively trust automated outputs, often using them as a cognitive shortcut to reduce mental effort (mosier2018human; lyell2017automation). This can lead to errors of commission (accepting incorrect AI-generated information) and errors of omission (failing to notice problems that the AI missed) (ABDELWANIS2024460).
On the other hand, algorithm aversion describes a tendency to prefer human judgment over algorithmic judgment, even in cases where the algorithm is demonstrably better or equivalent (Sunstein_Gaffe_2025). This aversion is often heightened in high-stakes, subjective, or ethically charged tasks, such as medical diagnosis (filiz2023extent). It can be driven by a fundamental desire for human agency, a negative emotional reaction to being judged by a machine, and a belief that human experts possess unique, ineffable knowledge that algorithms cannot capture.
Our study’s finding—that authors edit AI-attributed abstracts significantly more heavily—can be interpreted as a manifestation of algorithm aversion. The explicit knowledge of the AI source appears to trigger heightened scrutiny and a greater perceived need for human intervention. Furthermore, our study advances the literature by shifting the focus from perceptual measures of AI-generated text such as self-reported “credibility” or “trust” (LERMANNHENESTROSA2023107445; MaC3019) to concrete, measurable behavior in terms of amount of edits made to the text. The finding that attribution significantly impacts the number of edits demonstrates that this cognitive bias translates into tangible, effortful action on the part of the expert.

### v2

@@ Abstract.
Large Language Models have seen expanding application across domains, yet their effectiveness as assistive tools for scientific writing—an endeavor requiring precision, multimodal synthesis, and domain expertise—remains insufficiently understood. We examine the potential of LLMs to support domain experts in scientific writing, with a focus on abstract composition. We design an incentivized randomized controlled trial with a hypothetical conference setup where participants with relevant expertise are split into an author and reviewer pool. Inspired by methods in behavioral science, our novel incentive structure encourages authors to edit the provided abstracts to an acceptable quality for a peer-reviewed submission. Our 2 x 2 between-subject design expands into two dimensions: the implicit source of the provided abstract and the disclosure of it. We find authors make most edits when editing human-written abstracts compared to AI-generated abstracts without source attribution, often guided by higher perceived readability in AI generation. Upon disclosure of source information, the volume of edits converges in both source treatments. Reviewer decisions remain unaffected by the source of the abstract, but bear a significant correlation with the number of edits made. Careful stylistic edits, especially in the case of AI-generated abstracts, in the presence of source information, improve the chance of acceptance. We find that AI-generated abstracts hold potential to reach comparable levels of acceptability to human-written ones with minimal revision, and that perceptions of AI authorship, rather than objective quality, drive much of the observed editing behavior. Our findings reverberate the significance of source disclosure in collaborative scientific writing.
@@
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789140isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Telling people about research is as important as doing it. Scientific publishing is an integral part of science, meant to disseminate research findings, foster collaboration, encourage reproducibility, and ensure that scientific knowledge is accessible and built upon over time (Liang et al., 2024). Bad writing can and often does prevent or delay the publication of good science (Gastel and Day, 2022). Researchers and scientists today, however, are often not formally trained to write, and generally learn by imitating the styles of their advisors and other authors (Lindsay, 2020). Large language models (LLMs), with their writing abilities, offer a promising venue to assist researchers in both conducting and writing about research. However, prior work on thorough assessment of their capabilities in open ended scientific writing tasks remains limited.
@@
Even so, the use of language models continues to grow in this area. Many tools and platforms termed “AI co-scientist” have emerged in the past year (Gottweis et al., 2025; Beel et al., 2025). They are marketed to be capable of generating, executing, and writing papers, with some of them getting accepted in academic workshops organized by peer-reviewed conferences (Yamada et al., 2025). Popular scientific writing platforms such as Overleaf now offer writing assistant plugins. [^1] https://www.overleaf.com/learn/how-to/Writefull_integration However, anecdotal examples of both seemingly AI-generated papers (Cybernews, 2023; DeGeurin, 2024) and peer reviews (Oransky and Marcus, 2024) have inspired criticisms (Liang et al., 2024). In addition, across disciplines, recent studies have raised concerns over the declining quality of writing in scientific publications in several disciplines, attributing it to the rise of LLM-generated content (Matsui, 2024; Lemire, 2024; Geng and Trotta, 2024). These studies, however, have largely focused on corpus-level trends on unreviewed manuscripts and are conducted without consulting or involving writers such as the authors of said publications.
To raise awareness and prevent misuse by bad actors or unintentional use by over-reliant writers, there is an urgent need to conduct a systematic evaluation of the writing abilities of the current state-of-the-art models in specialized domains. Furthermore, despite current limitations, AI assistance holds tremendous potential in accelerating future scientific progress (Sourati and Evans, 2023; Nathani et al., 2025; Majumder et al., 2024a). Towards understanding scientific writing capabilities, we seek to answer the following research question in this work: maintaining the standard for scientific writing quality, to what extent do researchers rely on model-generated content for academic publishing?
@@
We simulate an abstract editing and reviewing process mimicking the process of scientific publication in a laboratory setting using the methodology of incentivized behavioral experiments (Azrieli et al., 2018) to encourage participants to perform the task to the best of their ability. Authors edit the abstracts, and we quantify author reliance based on the extent of their modifications to make them submission-ready for a hypothetical conference publication. The task is conducted under two source information conditions: (1) authors are informed that the original abstract provided to them was AI-generated, and (2) authors are not informed that the abstract was assembled by generative AI. To contextualize this behavior with a more realistic workflow, we also provide the authors with human-written abstracts under the same source information conditions. By considering abstracts from different sources (human experts and LLMs), we assess the reception of LLM-generated abstracts by evaluating correlations between acceptance rates, edit frequency, and edit types.
@@
The edited abstracts are then evaluated by a group of (incentivized) reviewers in a double-blind fashion, who vote to accept or reject the submitted abstract. This design allows us to examine how authors’ reliance varies across source types and source information conditions, and how expert reviewers perceive the quality and credibility of abstracts produced through generative AI workflows. These experiments result in a collection of model-generated and human-edited texts, where we capture keystroke-level edits. From this rich data of inputs, abstracts, and edits, we conduct several evaluations and analyses. Our main evaluation is based on the correlation between the rate of editing and the final rating an abstract receives, contrasting the controls and different treatments. We consider and measure the impact of variables like writer expertise (undergraduate students, graduate students, experienced researchers), and demographics and social factors (gender, perception of AI). In addition, we quantitatively (using linguistic tools) and qualitatively (via interviews with participants) characterize different kinds of edits authors make to understand the differences between edits made to AI- and human-authored abstracts.
@@
Across our experiments, we find that both the implicit origin of abstract and the disclosure of it systematically shape editing behavior and outcomes. Without source disclosure, authors make fewer edits to AI-generated abstracts, perceiving them as more readable, though PhD-level authors edited AI text more extensively. Despite these differences, reviewer evaluations showed no differences in accepting abstracts edited under no disclosure conditions. Disclosure of source primarily triggers social adjustments: authors reduced edits for human-written text when the source was revealed, whereas disclosure had only minor effects on AI-generated abstracts, with small impacts on perceived readability and confidence.
Our quantitative analysis of stylistic metrics reveals that edits to AI-generated abstracts are to improve cohesion, reduce nominalizations, and produce informative sentences. On the contrary, authors tend to lengthen the openings of the human-written abstracts. Similarly, disclosure of source alone has a minimal direct effect on the writing style. Finally, thematic analysis of our interviews with authors confirms that authors adopt different strategies depending on the origin of the abstracts: they simplify AI-generated text while restructuring human-authored abstracts for emphasis. Disclosure promotes a sense of accountability for both human- and AI-authored texts. Together, through a comprehensive large-scale experiment, we first highlight the varying cognitive editorial strategies emerging from writing that involves both human and AI-generated scientific writing. [^2] Our experimental data, task instructions, and example of a task are provided as supplementary.
@@ 2.1. AI Systems for Automating Scientific Research
Early works on AI-assistance in scientific research largely focused on surface-level corrections and sentence-level improvements in writing (Ito et al., 2019; Bezerra et al., 2021) with some attempts at drafting entire paper structure (Wang et al., 2019). Recent work has started to apply LLMs to the entire research cycle, including hypothesis discovery (Zhou et al., 2024; Zhang et al., 2025; Majumder et al., 2024b; Abdel-Rehim et al., 2025), experiment planning and execution (Li et al., 2025; Binbas, 2025; Gui and Kim, 2025), scientific writing (Kanna, 2024; Liebling et al., 2025; Kousha and Thelwall, 2025), and writing reviews (Silva et al., 2025). Within the writing stage, some works have focused on specific subtasks like metaphor generation (Kim et al., 2023), figure caption generation (Kim et al., 2025), citation generation (Xing et al., 2020; Fierro et al., 2024; Jung et al., 2022), and related work generation (Funkquist et al., 2022; Li and Ouyang, 2024). Such tasks are, typically, accompanied by reference answers, making evaluation relatively straightforward. More recent work has started focusing on generating entire papers without any human intervention (Lu et al., 2024). In this work, we adopt a more realistic setup of humans relying on an AI-assistant to generate drafts, which they then edit (Kanna, 2024). We also use a principled method of human evaluation in the form peer review conducted via incentivized experiments.
A wide range of commercial [^3] https://www.gatsbi.com/, https://anara.com/blog/ai-research-tools as well as open-source tools now offer services for every stage of the writing process, from literature review and synthesis to full manuscript drafting, editing, and refinement (Shao et al., 2025). Specific features, such as automated abstract and title generators, are also becoming commonplace. [^4] https://www.writefull.com/ The widespread availability and marketing of these tools signify a de facto adoption that necessitates rigorous empirical evaluation of their effectiveness, their impact on the quality of scientific output, and the human behaviors they engender, which our study addresses. In fact, large-scale corpus analyses suggest that AI assistance in scientific writing is already widespread (Kobak et al., 2025) in certain domains. Recent studies emphasize the need to understand how human researchers interact with AI-generated drafts, how much agency they retain, and how these systems reshape scientific communication norms (Reza et al., 2025). Our work contributes to this line of inquiry by focusing on human reliance and editorial behavior when interacting with AI-generated scientific text under realistic publication incentives.
@@ 2.2. Human AI Collaboration and Co-writing
The majority of prior research on human-AI co-writing has focused on creative writing (Singh et al., 2025). Using LLMs, researchers have developed and evaluated systems supporting story writing (Yuan et al., 2022), playwriting (Grigis and De Angeli, 2024), and character development (Park et al., 2025). Some works also explore higher-level writing tasks such as prewriting (Wan et al., 2024), and generating perspective-specific feedback (Rashkin et al., 2025). Specialized creative applications have emerged for tasks including metaphor generation (van Heerden and Bas, 2024), collaborative storytelling (Wang et al., 2024), and personal diary writing (Kim et al., 2024). User studies reveal complex dynamics in how writers interact with and perceive AI writing assistance. At a system interaction level, the design of AI suggestions significantly impacts user behavior and output: sentence-level suggestions promote original content creation, while paragraph-level suggestions improve efficiency (Fu et al., 2023; Dhillon et al., 2024). Writers’ engagement with AI assistance is also influenced by their personal values and goals (Biermann et al., 2022; Guo et al., 2025). Writers show varying receptivity to AI support based on their confidence levels, demonstrating higher acceptance in areas where they lack expertise (Al-Bukhrani et al., 2025; Wang and Lu, 2025), and their desires for support are closely tied to their perception of support actors and personal values (Hwang et al., 2025; Guo et al., 2024). Professional writers note persistent challenges with AI systems’ ability to maintain consistent style and voice (Chakrabarty et al., 2024a).
Moreover, this human-AI writing relationship raises important concerns. Studies reveal that biased AI models can influence not only the resulting text but also users’ own opinions (Jakesch et al., 2023). Prior research has identified biases, quality issues, lack of structure, and superficiality in AI-generated text, especially in domains like fiction and satire (Chakrabarty et al., 2024a, b; Ippolito et al., 2022; Marco et al., 2024; Mirowski et al., 2023, 2024). Researchers have also found LLM-generated content to be overly positive and lacking in nuance and have highlighted the need for more diverse and representative training data to mitigate these shortcomings (Tian et al., 2024). Our work adapts evaluation practices from prior studies in co-writing by focusing on the quality of AI-generated text using the amount of edits made to it by a human co-writer. Related to our work, experimental studies demonstrate that higher-level AI scaffolding can improve writing quality and efficiency, particularly for less experienced writers, while simultaneously raising concerns around diminished authorship and ownership (Dhillon et al., 2024). In parallel, large-scale reviews of human–AI co-writing systems emphasize that writers desire different forms of AI support at different stages of the writing process, such as ideation, drafting, and revision (Reza et al., 2025). This underscores the importance of evaluating AI assistance within realistic workflows rather than isolated generation tasks.
@@ 2.3. Cognitive Biases in Human-AI interaction
The inability of humans to reliably detect AI-generated text makes their beliefs about a text’s origin a powerful psychological variable, opening the door for cognitive biases to influence their judgment and behavior (Jakesch et al., 2022; Henestrosa and Kimmerle, 2024; Fiedler and Döpke, 2025). A rich literature in HCI and psychology documents two opposing, context-dependent human tendencies when interacting with automated systems (Parasuraman et al., 2000; Lee and See, 2004; Dzindolet et al., 2003). On one hand, automation bias describes the tendency to over-rely on or excessively trust automated outputs, often using them as a cognitive shortcut to reduce mental effort (Mosier and Skitka, 2018; Lyell and Coiera, 2017). This can lead to errors of commission (accepting incorrect AI-generated information) and errors of omission (failing to notice problems that the AI missed) (Abdelwanis et al., 2024).
On the other hand, algorithm aversion describes a tendency to prefer human judgment over algorithmic judgment, even in cases where the algorithm is demonstrably better or equivalent (Sunstein and Gaffe, 2025). This aversion is often heightened in high-stakes, subjective, or ethically charged tasks, such as medical diagnosis (Filiz et al., 2023). It can be driven by a fundamental desire for human agency, a negative emotional reaction to being judged by a machine, and a belief that human experts possess unique, ineffable knowledge that algorithms cannot capture.
Our study advances the literature by shifting the focus from perceptual measures of AI-generated text such as self-reported “credibility” or “trust” (Lermann Henestrosa et al., 2023; Graefe and Bohlken, 2020) to concrete, measurable behavior in terms of amount of edits made to the text. The finding that attribution significantly impacts the number of edits demonstrates that this cognitive bias translates into tangible, effortful action on the part of the expert, also mirrored by recent empirical work (Zhu et al., 2025). Controlled experiments in German theses by Fiedler and Döpke (2025) show that when identical texts are labeled as AI-generated versus human-written, evaluators systematically favor the human-labeled versions, even when the labels are incorrect, demonstrating a strong source-identity bias against AI authorship. Conversely, when source information is withheld, both expert and non-expert evaluators struggle to reliably distinguish AI-generated academic writing from human writing, often rating them as comparable in quality.